f(x) = 2x^2 - 1 and g(x) = 3x + 2
Find each value.
f(g(-2))
g(f(sqrt2))

Answer by stanbon(48564)

f(x) = 2x^2 - 1 and g(x) = 3x + 2
Find each value.
f(g(-2))=f(3*-2+2)=f(-4)=2(-4)^2-1=15
g(f(sqrt2))=g(2(sqrt2)^2-1)=g(3)=3*3+2=11
-----------
Cheers,
Stan H.